A restaurant reported the following current assets: cash P12,000; credit card receivables P1,800; accounts receivable P180; food inventory P4,400; and prepaid expenses P1,120. Current liabilities total P7,800. Answer the following:
a. Calculate the current ratio.
b. Calculate the quick ratio (acid-test ratio).

Problem 8.
A restaurant and beverage operation reported the following for the operating month of March. March has 23 operating days.

March 1 March 31 Cost of Sales
Food service inventory P 8,868 P 5,740 P 36,520

For the month of March, calculate the food inventory turnover food ratio and average period in days it takes for food inventory to turn over.
1
Expert's answer
2021-09-26T20:13:51-0400

Solution:

Problem 7:

a.). Current ratio = "\frac{Current\;Assets}{Current \;Liabilities}"


Current assets = Cash + Credit card receivables + Accounts receivable + Food inventory + Prepaid expenses

Current assets = 12,000 + 1,800 + 180 + 4,400 + 1,120 = 19,500

Current liabilities = 7,800


Current ratio = 19,500/7,800 = 2.5


Current ratio = 2.5

 

b.). Quick ratio (acid-test ratio) = "\frac{Cash + Credit\;card \;receivables + Accounts\;Receivable}{Current \;Liabilities}"


="\frac{12,000 + 1,800 + 180}{7,800} = \frac{13,980}{7,800} = 1.79"


Quick ratio (acid-test ratio) = 1.79

 

Problem 8:

The food inventory turnover food ratio = "\frac{Cost\;of\;goods\;sold}{Average\;Inventory}"

Cost of goods sold = 36,520


Average inventory = "\frac{(8,868 + 5740)}{2} = 7,304"


The food inventory turnover food ratio = "\frac{36,520}{7,304} = 5"


The food inventory turnover food ratio = 5

 

Average period in days it takes for food inventory to turn over = "\frac{Period}{The \; food \; inventory \; turnover \; ratio}" = "\frac{23}{5} = 4.6 \; days"